# Tailored Forecasting from Short Time Series via Meta-learning

### v1

@@ II.4 Simultaneous learning and generalization of logistic and Gauss iterated maps
II.5 Generalization in fully and partially observed Lorenz-63 systems
Figure 5: An example Lorenz-63 forecast using METAFORS
The test signal, of ${N_{test}=30}$ sequential points, is the segment of the true signal before the vertical dashed black line at ${t_{test}\equiv(N_{test}-1)\Delta t}$. The forecast starts at ${t_{test}}$ and yields a valid prediction time ${T_{valid}=340\Delta t}$, as defined by Eq. 1. The parameters of the $N_{L}=9$ long library signals are as shown in Fig. 6. They were selected randomly from the uniform distributions ${\omega_{t}\in U[0.75,1.25]}$ and ${v_{1}\in U[7.5,12.5]}$, as in Section IV.3.2. The Lorenz parameters of the test system are ${(\omega_{t},v_{1})\approx(1.25,10.83)}$.
While we believe that larger applications like weather forecasting could benefit from the approaches used in METAFORS, here we employ a simplified but widely studied model of atmospheric convection, the Lorenz-63 equationsLorenz (1963), as another test bed for METAFORS. First, we briefly demonstrate that METAFORS’ forecasts capture the climate of test systems with unseen Lorenz-like dynamics, and then we use METAFORS to predict the short-term evolution of such systems in a few distinct scenarios. Of particular note, we show that METAFORS can still generalize and cold-start the forecaster when only partially-observed states of the training and test systems are available. Partial-state forecasting is required in many applications – especially those where the system of interest is high-dimensional and only a few variables can be feasibly measured, such as in weather-forecasting and epidemiology.
@@ II.5 Generalization in fully and partially observed Lorenz-63 systems
Each set of values of the parameters $\omega_{t}$, $v_{1}$, $v_{2}$, and $v_{3}$ define a unique dynamical system. For our experiments, every long library signal and short test signal is a segment of the attractor of a Lorenz system with different values of the parameter $v_{1}$ and the time-scale, $\omega_{t}$. We hold ${v_{2}=28}$ and ${v_{3}=8/3}$ fixed. The factor $\omega_{t}$ does not appear in the original formulation of the equations, but allows for changes to the time-scale of chaotic oscillations. Increasing $\omega_{t}$ causes a system to evolve more rapidly with time.
For all values of the Lorenz parameters that we explore here, the equations exhibit chaotic dynamics: any two trajectories will become decorrelated after a finite time, regardless of how close their initial conditions are. We can quantify the time-scale over which a system’s underlying chaotic dynamics make prediction impossible using Lyapunov timesTél and Gruiz (2006). A Lyapunov time is the time required for the distance between two trajectories that are initially close together to increase by a factor of Euler’s number, $e$. A system’s Lyapunov time scales proportionally to $\omega_{t}$. We therefore measure periods of time in units of our constant time step, ${\Delta t=0.01}$, to report results. For context, however, we note that the Lyapunov time of the Lorenz system with standard parameter values ${\omega_{t}=1}$, ${v_{1}=10}$, ${v_{2}=28}$, and ${v_{3}=8/3}$ is $\tau_{Lyap}\approx 1.104\approx 110\Delta t$. Because we use a constant time step, by varying $\omega_{t}$ we vary how finely the data are sampled.
@@
In our experiments with fully-observed Lorenz systems, we train the forecaster RC to take as input and predict as output the full system state ${(x_{1},x_{2},x_{3})}$ at each instant. For our experiments with partially-observed Lorenz systems, the forecaster is trained on and predicts the $x_{3}$-component only. Full details of the experimental parameters defining METAFORS’ library and training scheme for our experiments with Lorenz systems are given in Section IV.3.2. An example forecast of a fully-observed Lorenz system using METAFORS is shown in Fig. 5. Note that the test signal starts at time ${t=0}$ and the first predicted data-point is at $t=t_{test}$. For clarity, we follow this convention always. The valid time is measured as in Eq. 1.
II.5.1 METAFORS captures both the climate and short-term evolution of unseen Lorenz systems
Figure 6: METAFORS generalizes the forecaster to Lorenz-like dynamics unseen in the library.
The Lorenz parameters of the long library signals, marked by black dots, were randomly selected from the uniform distributions ${\omega_{t}\in U[0.75,1.25]}$ and ${v_{1}\in U[7.5,12.5]}$, as in Section IV.3.2. The library and test signals contain observations of the full state of the Lorenz system. (A) At each grid point, we calculate the mean autonomous one-step error, which measures climate reproduction, for each of $500$ forecasts of duration $3000\Delta t$, and then plot the median of these means. (B) The median valid times, which measure short-term prediction accuracy, for the same set of forecasts. In (A1) and (B1), we use the forecaster of the library member closest in ${(\omega_{t},v_{1})}$-parameter space to the test signal. In (A2) and (B2), we triangulate the parameter values ${(\omega_{t}^{test},v_{1}^{test})}$ of each test system with respect to the Lorenz parameter values of the library members. If the test parameters lie inside the convex hull of the long library signals, we calculate the output layer of the forecaster RC by interpolation of the three closest library members. Otherwise, we use the nearest library forecaster. The test length, ${N_{test}=200}$, is longer than the memory of the forecaster, so that the cold-start vector learned by METAFORS offers no advantage.
@@ II.5.1 METAFORS captures both the climate and short-term evolution of unseen Lorenz systems
In Fig. 6, we present results for a library constructed from nine Lorenz systems with different parameter values, illustrating how the long-term climate replication (A) and short-term forecast accuracy (B), measured by the autonomous one-step error and valid prediction time, respectively, vary with the test parameters $v_{1}$ and $\omega_{t}$. We compare METAFORS’ generalization ability to a few simple baseline approaches. First, for the Nearest Library Forecaster, we rescale the dynamical parameters $\omega_{t}$ and $v_{1}$ used to generate the long library signals such that they span a unit interval along both axes. Then, for each test signal, we make a prediction using the forecaster RC of the long library signal whose re-scaled dynamical parameters are nearest to those of the test system. We emphasize that in typical applications we do not know the dynamical parameters associated with either the short test signals or the long library signals, so this comparison is typically infeasible. For the Interpolated Forecaster (Section IV.4.3), we perform linear interpolation of the forecaster model parameters if the test system lies within the convex hull of the library. Otherwise, we use the forecaster RC of the nearest library member. We choose this interpolation/extrapolation method because it is applicable to high-dimensional spaces and assumes no underlying structure in the distribution of the library members, but more sophisticated schemes may offer better performance. As with the nearest library forecaster, we note that interpolation of the model parameters is typically infeasible. We include it here as a comparison to a scheme that relies on additional information (i.e., the typically unknown dynamical parameters). Since no cold-start vector is learned in either of these methods, or in our simple multi-task learning approach, we obtain forecasts by synchronizing the forecaster to the test signal from a zero-vector internal state, $\bm{r}(0)=\bm{0}$, before prediction begins at ${t=t_{test}}$.
@@
Fig. 6(A1) and (B1) illustrate the brittleness of the RC forecasting models in our library. Each forecaster model in the library provides strong climate replication (A1) and accurate short-term forecasts (B1) only if the test signal parameters are very close to those of the system on which it was trained. Panel (A4), on the other hand, demonstrates that our basic multi-task learning approach to training the forecaster fails to capture the climate of unseen test systems even when their dynamical parameters are quite close to those used in the library. The multi-task learning approach is slightly more useful for short-term forecasting (B4), but struggles in a way that is common to multi-task learning methods: by training to improve performance on the library members generically, it does not forecast any single system accurately. We show in Fig. 10 that a larger (i.e., more powerful) multi-task forecaster RC does not close the performance gap between the basic multi-task learning method and METAFORS. Panels (A3) and (B3) demonstrate that METAFORS facilitates the generalization that is required for this problem. Over a wide range of test signal dynamics, METAFORS offers lower autonomous one-step errors (A3) and higher valid prediction times (B3) than the other methods. In particular, METAFORS offers longer valid prediction times than does the parameter-aware Interpolated Forecaster (B2) method unless the test dynamics are very similar to the dynamics of one of the library members.
@@
Fig. 11 explores further how the library structure (i.e., the length and number of library signals) affects METAFORS’ performance.
@@
II.5.2 With very short test signals, METAFORS’ cold-starting is essential
Figure 7: METAFORS forecasts partially-observed Lorenz systems from very limited data. Using a library comprising $N_{L}=9$ long signals with Lorenz parameters indicated by the black dots in Fig. 6, we calculate mean valid times over $625$ test signals arranged in a $25\times 25$ rectangular grid spanning the space defined by ${0.7\leq\omega_{t}\leq 1.3}$ and ${7\leq v_{1}\leq 13}$. The training and test signals contain only the $x_{3}$ Lorenz variable. Error bars denote the standard error of the mean. For METAFORS, No Cold-start (blue), the signal mapper RC learns model parameters but no cold-start vector for the forecaster. For that method and all others except METAFORS, we synchronize the forecaster to the test signal from a zero-vector internal state, ${\bm{r}(0)=\bm{0}}$, and then predict in autonomous/closed-loop mode (Fig. 2B) from the end of the test signal.
@@ II.5.2 With very short test signals, METAFORS’ cold-starting is essential
We plot mean valid prediction time against test signal length for METAFORS and a few baseline methods in Fig. 7. Here, both the training and test signals contain only partially-observed Lorenz states (the $x_{3}$-component only). We present similar results with fully-observed Lorenz systems in Fig. 12. METAFORS’ ability to cold-start forecasts strikingly increases valid prediction times when the test signals are short. With test signals of length $t_{test}=19\Delta t$ (${N_{test}=20}$), for instance, METAFORS’ mean valid prediction time, ${T_{valid}\approx 139\Delta t}$, is just over seven times the length of the test signals. All other methods synchronize the forecaster RC to the test signal from a zero-vector internal state ${\bm{r}(0)=\bm{0}}$ and thus cannot offer useful forecasts until the duration of the test signal is comparable to that of the forecaster’s memory. When the test signals are long enough that cold-starting is not required, METAFORS’ valid prediction time still settles at a higher plateau value than even our parameter-aware Interpolated Forecaster method, the best performing of the baseline approaches considered. The poor performance of forecaster RCs trained directly on the test signals (Train on Test Signal), in contrast, highlights that even these longest test signals are significantly shorter than is required to train a forecaster RC accurately.
II.5.3 METAFORS is useful for cold-starting even when generalization is not required
Figure 8: METAFORS is useful for cold-starting even when generalization is not required. Using a library comprising ${N_{L}=1}$ long time series, this single training signal and all $625$ test signals are segments of the attractor for the standard Lorenz system, ${\omega_{t}=1}$ and ${v_{1}=10}$, with different initial conditions. Since there is only one library member, the signal mapper learns only a cold-start vector, ${\bm{r}(0)}$, for the forecaster. For all methods, we train the forecaster parameters on the library signal directly. The training and test signals are partially-observed, containing only the $x_{3}$ Lorenz variable. Error bars denote the standard error of the mean.
@@ II.5.3 METAFORS is useful for cold-starting even when generalization is not required
While we typically do not expect ML forecasting models trained on data from one dynamical system to generalize well to test systems with different dynamics, we do expect that a forecaster should offer useful predictions for test signals that exhibit identical dynamics to those of its training system. ML forecasters with memory, however, struggle in this scenario unless the test signal is sufficiently long to initialize their memory.
@@
In Fig. 8, we train METAFORS on a single long library signal with standard Lorenz parameters ${\omega_{t}=1}$ and ${v_{1}=10}$ and test it on short signals with the same dynamics starting from $625$ different initial conditions. Since there is only one training time series, the signal mapper only has to map short signals to cold-start vectors, with the forecaster’s model parameters determined by training on the sole long library signal. In both the partially-observed (Fig. 8) and fully-observed (Fig. 12(B)) forecasting tasks, METAFORS’ signal mapper enables simple cold-starting of the forecaster. We emphasize that the signal mapper requires no more training data than is traditionally required for forecasting of stationary dynamics; it is trained from the same data as the forecaster in order to cold-start predictions at new points in phase space.
The utility of METAFORS’ cold-starting in this simplified setting is highlighted by comparison to the traditional approaches that we include as baseline methods in Fig. 8. For the method labeled ${\bm{r}(0)=\bm{0}}$, we synchronize the forecaster to each test signal starting from a zero-vector internal state before prediction begins. This is a typical approach for initializing a memory-based forecaster. In Backward Extrapolation as a Constant, we extrapolate the test signal backwards from its initial value, ${\bm{s}_{test}(0)}$, as a constant for two-hundred time steps and then synchronize the forecaster to this extrapolated signal
starting from ${\bm{r}(-200\Delta t)}=\bm{0}$. In Training Data Search, we search the training series for the segment, $s_{match}$, that minimizes the root-mean-square error between the test signal and itself. Denoting by $\bm{r}_{train}(t_{match})$ the constructed cold-start vector at the time step of the training signal at which $s_{match}$ begins, we then synchronize the forecaster to the test signal starting from the initial state $\bm{r}(0)=\bm{r}_{train}(t_{match})$. This method of searching the library for the closest match to the test signal offers considerable improvement over the two other traditional approaches that we highlight, but it requires a computationally expensive search each time we wish to make a new prediction. Its performance may also depend on how well the training signal covers its associated attractor. METAFORS, on the other hand, can learn a cold-start vector cheaply for each test signal – no retraining of the signal mapper is required.
We make two brief technical observations. METAFORS’ peak valid prediction time in Fig. 7 is significantly lower than it is in Fig. 8, where we do not require generalization to new dynamics. This discrepancy is related to training regularization. In Fig. 13, we show that in this case, with data from dynamical systems governed by the Lorenz-63 equations, our RC implementation of METAFORS benefits from using a high regularization strength ($10^{-6}$) for the linear regression step of the forecaster RC’s training. When training a forecaster RC optimally for prediction of test systems with identical dynamics to the training system, a lower regularization strength (${10^{-13}}$) provides better performance. This low regularization strength allows for an excellent fit to the training dynamics, and for higher valid prediction times than are achievable when generalization is required (Fig. 7). When generalization to new dynamics is our goal, a higher regularization prevents over-fitting to the training system but limits peak performance.
@@
In both the partially-observed (Figs. 7 and 8) and fully-observed (Fig. 12) cases, METAFORS offers useful forecasts with test signals consisting of only one data point (the point in phase space from which the forecast should start). While cold-starting from such a short test signal is noteworthy, these library and test signals contain no observational noise. In Fig. 14, we demonstrate that our RC implementation of METAFORS is robust to small amounts of observational noise in the test signals.
III Discussion
@@
This work introduces Meta-learning for Tailored Forecasting from Related Time Series (METAFORS), a framework designed to address significant challenges in forecasting dynamical systems from limited observations. By leveraging a library of machine learning models trained on related systems (for which extensive data are available), METAFORS demonstrates the ability to construct and initialize forecasting models tailored to test systems with limited observations, without requiring any explicit knowledge of the test or training systems’ underlying dynamics. Our study demonstrates the capabilities of METAFORS in multiple nonlinear systems, demonstrating its robustness, versatility, and potential applicability to real-world systems.
METAFORS addresses two critical challenges commonly faced by memory-based machine learning (ML) forecasters: brittleness and the requirement for substantial system-specific training data. By constructing a “signal mapper” to utilize a library of individually trained ML models, METAFORS identifies suitable forecasting model parameters and cold-start vectors from short observation signals. The results presented here highlight METAFORS’ unique advantages:
  1. 1.
    Generalization Across Systems: METAFORS successfully forecasts the behavior of test systems with dynamics that are related to, yet substantially different from, those in the training library. It achieves this generalization without contextual awareness or knowledge of governing dynamical parameters, as evidenced by its ability to replicate the logistic map’s bifurcation diagram, simultaneously learn and generalize the logistic and Gauss iterated maps, and generalize to Lorenz dynamics generated with dynamical parameters that differ substantially from those used to construct the training data.
  2. 2.
    Cold-Starting Forecasts: The ability to cold-start memory-based models is crucial when test signals are very short. METAFORS consistently outperforms our baseline methods, even enabling useful forecasts with minimal initialization data. This capability is particularly important for systems where data collection is constrained, as is the case in many biomedical, epidemiological, and ecological applications.
  3. 3.
    Flexibility in Model Architecture: METAFORS is not tied to specific ML architectures or training schemes. In our examples, we use reservoir computing for both the forecaster and signal mapper, but the principles of METAFORS are general and can accommodate other ML implementations for the forecaster or signal mapper.
  4. 4.
    Ability to Simultaneously Capture Short-term Evolution and Long-term Behavior: METAFORS is able to simultaneously predict both short-term evolution and long-term statistics (climate) of dynamical systems, making it applicable to a wide range of forecasting scenarios.
  5. 5.
    Accuracy and Efficiency when Implemented with Reservoir Computing: While METAFORS is framework-agnostic, our implementation with reservoir computers (RCs) leverages their appealing features – simplicity, computational efficiency, and low data and initialization requirements – while also extending the utility of traditional reservoir computing to tackle more challenging forecasting problems that require generalization.
Previous work has demonstrated that memory-based ML models with input channels for explicit context-indicators can capture the climate of unseen systems whose behaviors are qualitatively different but whose dynamics are governed by equations of the same functional form Patel et al. (2021); Kong et al. (2021); Köglmayr and Räth (2024); Panahi and Lai (2024); Kong, Brewer, and Lai (2024). By contrast, METAFORS achieves this generalization with no such contextual awareness. Instead, METAFORS’ signal mapper learns the context directly from short cue signals and then both constructs and initializes a tailored forecasting model. Since METAFORS constructs models from observational data alone, it requires little domain-specific knowledge. In Section II.4, we exploit this versatility and train METAFORS to learn, simultaneously and without context, dynamical systems that are governed by two distinct functional forms (logistic and Gauss iterated maps). Parameter-aware generalization schemes cannot readily be extended to this setting without adding additional contextual indicators about the type of system from which training and test signals are drawn.
@@
While METAFORS’ data-driven modeling approach is an important strength, the integration of METAFORS with hybrid forecasting architectures—where knowledge-based models are coupled with data-driven components—could further enhance its utility. Exploring such integrations could enable METAFORS to address scenarios where partial knowledge of the system dynamics is available, combining the strengths of data-driven and knowledge-based approaches.
Although the framework of METAFORS is not tied to a specific ML architecture, it is noteworthy that the reservoir computing implementation of METAFORS we develop here not only inherits the simplicity, versatility, and computational efficiency of RCs generally (with training achieved by simple linear regression), but constitutes an important extension of reservoir computing itself. Yan et al.Yan et al. (2024) recently identified generalizable reservoir computing as a key to facilitating industrial use of RCs. In tempering the requirements of identical training and test dynamics (‘brittleness’), lengthy training series (‘data-intensity’), and warm-up signals to initialize predictions, which all constrain RCs in applications beyond forecasting, METAFORS constitutes an important step towards this goal.
We are aware of a small number of related approaches to forecasting dynamical systems from limited data. Two recent papers have introduced methods to cold-start RCsGrigoryeva et al. (2024) and LSTMsKemeth et al. (2021). However, these studies highlight the importance of cold-starting only in the context of short-term forecasting and do not address scenarios in which the test and training signals differ in both their dynamics and their initial conditions. Here, we show that a framework to generalize forecasting models more broadly across systems with varied dynamics readily simplifies for this more limited purpose. Similarly, other meta-learning frameworks for dynamical systems are constrained to particular implementations or applications, either because they cannot cold-start forecasting models with memory or because they depend on specific ML architectures. Some, for instance, are formulated explicitly using convolutional neural networks to capture spatiotemporal dynamicsWang, Walters, and Yu (2022) or rely on data-compressing ML architectures like autoencoders to perform the meta-learning taskCanaday, Pomerance, and Girvan (2021); Kirchmeyer et al. (2022); Panahi et al. (2025). METAFORS’ strength is its versatility: it both generalizes and cold-starts the forecasting model; it simultaneously predicts the test system’s short-term evolution and long-term climate; and it is flexible with respect to the architectures of both the forecaster and the signal mapper.
While METAFORS demonstrates significant promise, it also presents limitations that warrant further investigation. The current implementation requires uniformly sampled, sequential data for training and forecasting. Other implementations or extensions of METAFORS to handle irregularly sampled data or multiscale dynamics would broaden its applicability.
Finally, while we have demonstrated the utility of METAFORS in relatively simple nonlinear systems, its scalability to high-dimensional datasets and real-world applications remains an open question. Incorporating unsupervised learning techniques, such as autoencoders, to learn interpretable, low-dimensional representations of dynamics could improve METAFORS’ scalability and facilitate its application to more complex systems.

### v2

@@ Abstract
II.4Generalization in fully and partially observed Lorenz-63 systems
While we believe that larger applications like weather forecasting could benefit from the approaches used in METAFORS, here we employ a simplified but widely studied model of atmospheric convection, the Lorenz-63 equationsLorenz (1963), as another testbed for METAFORS. First, we briefly demonstrate that METAFORS’ forecasts capture the climate of test systems with Lorenz dynamics unseen in training, and then we use METAFORS to predict the short-term evolution of such systems in a few distinct scenarios. Of particular note, we show that METAFORS can still generalize and cold start the forecaster when only partially-observed states of the training and test systems are available. Partial-state forecasting is required in many applications – especially those where the system of interest is high-dimensional and only a few variables can be feasibly measured, such as in weather-forecasting and epidemiology.
@@
Each set of values of the parameters $\omega_{t}$, $v_{1}$, $v_{2}$, and $v_{3}$ define a unique dynamical system. For our experiments, the long library signals correspond to segments of the attractors of chaotic Lorenz systems with different values of the parameter $v_{1}$ and the time-scale, $\omega_{t}$. We hold ${v_{2}=28}$ and ${v_{3}=8/3}$ fixed. The factor $\omega_{t}$ does not appear in the original formulation of the equations, but presents the additional challenge of a varying time scale for the dynamics. Due to this variation, we present results in units of reservoir time steps (${\Delta t=0.01}$ in units of the differential equations) rather than in units of a characteristic dynamical time scale such as the Lyapunov time, which differs across systems. (The Lyapunov time is the typical amount of time required for the distance between two trajectories that are initially close together to increase by a factor of Euler’s number, $e$. It quantifies the time-scale over which a system’s chaotic dynamics make prediction impossibleTél and Gruiz (2006).) For context, however, we note that the Lyapunov time of the Lorenz system with standard parameter values ${\omega_{t}=1}$, ${v_{1}=10}$, ${v_{2}=28}$, and ${v_{3}=8/3}$ is $\tau_{Lyap}\approx 1.104\approx 110\Delta t$.
@@
Full details of the experimental parameters defining METAFORS’ library and training scheme for our experiments with Lorenz systems are given in Section˜.2.2. An example forecast of a fully-observed Lorenz system using METAFORS is shown in Fig. 5(A). Note that the test signal starts at time ${t=0}$ and the prediction starts at time ${t=t_{test}}$.
II.4.1Climate replication and forecasting for unseen Lorenz systems
Figure 5: METAFORS generalizes the forecaster to unseen fully-observed and partially-observed Lorenz sytems. (A) An example METAFORS forecast from a test signal of ${N_{test}=30}$ fully-observed Lorenz states with dynamical parameters ${(\omega_{t},v_{1})\approx(1.25,10.83)}$. The forecast, starting at ${t_{test}\equiv(N_{test}-1)\Delta t}$ (vertical dashed black line), yields a valid prediction time ${T_{valid}=340\Delta t}$. (B) The dynamical parameters used to generate the nine long library signals (black dots) for panels (A) to (C) were randomly selected from the uniform distributions ${\omega_{t}\in U[0.75,1.25]}$ and ${v_{1}\in U[7.5,12.5]}$. (B1) Median autonomous one-step error calculated over 500500500 full-state forecasts, each of duration $3000\Delta t$. (B2) Median valid times for the same set of forecasts. The test length, ${N_{test}=200}$, is longer than the memory of the forecaster, so that the cold-start vectors learned by METAFORS offer no advantage over the other methods. (C and D) Mean valid prediction time against test signal length with partially-observed training and test systems (only the Lorenz $x_{3}$-variable is measured) in two distinct scenarios. Error bars denote the standard error of the mean. (C) The dynamical parameters of the 625625625 test signals are arranged in a $\times 25$ 25 rectangular grid spanning ${0.7\leq\omega_{t}\leq 1.3}$ and ${7\leq v_{1}\leq 13}$. (D) The library contains only one long time series (${N_{L}=1}$). This single training time series and all 625625625 test signals have different initial conditions but the same dynamical parameters, ${\omega_{t}=1}$ and ${v_{1}=10}$. No generalization to new dynamics is required, so we train the forecaster parameters on the sole library signal directly (for both methods) and METAFORS’ signal mapper learns only a cold-start vector, ${\bm{r}(0)}$, for the forecaster.
@@
In Fig. 5(B), we present results for a library constructed from nine fully-observed Lorenz systems with different parameter values randomly chosen from the ranges ${0.7\leq\omega_{t}\leq 1.3}$ and ${7\leq v_{1}\leq 13}$, illustrating how the long-term climate replication (B1) and short-term forecast accuracy (B2), measured by the autonomous one-step error and valid prediction time, respectively, vary with the test parameters. We compare METAFORS’ generalization ability to a few simple baseline approaches. First, for the Nearest Library Forecaster, we rescale the dynamical parameters $\omega_{t}$ and $v_{1}$ used to generate the long library signals such that they span a unit interval along both axes. Then, for each test signal, we make a prediction using the forecaster RC of the long library signal whose re-scaled dynamical parameters are nearest to those of the test system. For the Interpolated Forecaster (Section˜.3.3), we perform linear interpolation of the forecaster model parameters if the test system lies within the convex hull of the library. Otherwise, we use the forecaster RC of the nearest library member. We emphasize that in typical applications we do not know the dynamical parameters associated with either the short test signals or the long library signals. So, both this method and Nearest Library Forecaster are typically infeasible. We include them for comparison to schemes that rely on additional information (i.e., the typically unknown dynamical parameters). Since no cold-start vector is learned in either of these methods, in our simple multi-task learning approach, or when training a forecaster on the test signal directly, we ‘zero start’ these models. That is, we obtain forecasts by synchronizing the forecaster to the test signal from a zero-vector internal state, $\bm{r}(0)=\bm{0}$, before prediction begins at ${t=t_{test}}$.
@@
Fig. 5(B1.1) and (B2.1) illustrate the brittleness of the RC forecasting models in our library. Each forecaster model in the library provides strong climate replication (B1.1) and accurate short-term forecasts (B2.1) only if the test signal parameters are very close to those of the system on which it was trained. Panel (B1.4), on the other hand, demonstrates that our basic multi-task learning approach to training the forecaster fails to capture the climate of unseen test systems even when their dynamical parameters are quite close to those used in the library. The multi-task learning approach is slightly more useful for short-term forecasting (B2.4), but struggles in a way that is common to multi-task learning methods: by training to improve performance on the library members generically, it does not forecast any single system accurately. We show in Fig. S2 that a larger (i.e., more powerful) multi-task forecaster RC does not close the performance gap between the basic multi-task learning method and METAFORS. Panels (B1.3) and (B2.3) demonstrate that METAFORS facilitates the generalization that is required for this problem. Over a wide range of test signal dynamics, METAFORS offers lower autonomous one-step errors (B1.3) and higher valid prediction times (B2.3) than the other methods. In particular, METAFORS offers longer valid prediction times than does the parameter-aware Interpolated Forecaster (B2.2) method unless the test dynamics are very similar to the dynamics of one of the library members.
@@
Fig. S3 explores further how the library structure (i.e., the length and number of library signals) affects METAFORS’ performance.
@@
II.4.2With very short test signals, proper cold starting is essential
@@
We plot mean valid prediction time against test signal length for METAFORS and a few baseline methods in Fig. 5(C). Here, both the training and test signals contain only partially-observed Lorenz states (the $x_{3}$-component only). We present similar results with fully-observed Lorenz systems in Fig. S5(A). METAFORS’ ability to cold start forecasts strikingly increases valid prediction times when the test signals are short. With test signals consisting of ${N_{test}=20}$ data points, for instance, METAFORS’ mean valid prediction time, ${T_{valid}\approx 139\Delta t}$, is just over seven times the length of the test signals. All other methods synchronize the forecaster RC to the test signal from a zero-vector internal state ${\bm{r}(0)=\bm{0}}$ (zero starting) and thus cannot offer comparably long valid times until the duration of the test signal is similar to that of the forecaster’s memory. When the test signals are long enough that cold starting is not required, METAFORS’ valid prediction time still settles at a higher plateau value than even our parameter-aware Interpolated Forecaster method, the best performing of the baseline approaches we consider. The comparably poor performance of forecaster RCs trained directly on the test signals (Train on Test Signal), in contrast, highlights that even these longest test signals are much shorter than is required to train an accurate RC forecaster.
II.4.3Effective cold starting when generalization isn’t required
@@
While we typically do not expect ML forecasting models trained on data from just one dynamical system to generalize well to test signals with different dynamics, we do expect that a forecaster should offer useful predictions for test signals that exhibit essentially the same dynamics as those of its training system. ML forecasters with memory, however, still struggle in this scenario unless the test signal is sufficiently long to initialize their memory.
@@
In Fig. 5(D), we train METAFORS on a single long library signal with standard Lorenz parameters ${\omega_{t}=1}$ and ${v_{1}=10}$ and test it on short signals with the same dynamics starting from 625625625 different initial conditions. Since there is just one training time series, the signal mapper only has to map short signals to cold-start vectors, with the forecaster’s model parameters determined by training on the sole long library signal. In both the partially-observed case (Fig. 5(D)), where we include only the $x_{3}$-component of the Lorenz system in the training signal and the test signals, and the fully-observed case (Fig. S5(B)), METAFORS’ signal mapper enables simple cold-starting of the forecaster. We emphasize that the signal mapper requires no more training data than is traditionally required for forecasting of stationary dynamics; it is trained from the same data as the forecaster in order to cold start predictions at new points in state space.
The utility of METAFORS’ cold starting in this simplified setting is highlighted by comparison to common elementary approaches to initializing a memory-based ML forecaster. We compare METAFORS’ performance to a number of alternative initialization methods in Fig. S4, but focus here (Fig. 5D) on one typical approach, zero starting.
@@
We make two brief technical observations. METAFORS’ peak valid prediction time in Fig. 5(C) is substantially lower than it is in Fig. 5(D), where we do not require generalization to new dynamics. This discrepancy is related in part to training regularization (Fig. S6). In both the partially-observed (Fig. 5C and D) and fully-observed (Fig. S5) cases, METAFORS offers useful forecasts with test signals of just one data point (the point in state space from which the forecast should start). While cold starting from such limited data is noteworthy, these test signals contain no noise. In Fig. S7, we demonstrate that our RC implementation of METAFORS is robust to small amounts of observational noise in the test signals.
BrickRedIIIDiscussion
@@
This paper introduces Meta-learning for Tailored Forecasting using Related Time Series (METAFORS), a framework designed to address key challenges in applying traditional ML approaches to forecasting dynamical systems, specifically their dependence on abundant system-specific training data and their brittleness when generalizing to related but distinct systems. By leveraging a library of machine learning models trained on related systems with ample data, METAFORS constructs and initializes tailored forecasting models for unseen systems using only limited observations and no additional contextual information. Our study exhibits METAFORS’ capabilities in multiple families of nonlinear systems, demonstrating its robustness, versatility, and potential applicability to real-world systems.
METAFORS offers several key features: (1) Generalization Across Systems: It forecasts test systems with dynamics that are related to but substantially different from those in the training library, without requiring contextual awareness. (2) Cold Starting Forecasts: It enables memory-based models to generate accurate forecasts from minimal initialization data, outperforming baselines – an essential capability for data-limited applications. (3) Flexibility in Model Architecture: METAFORS is not restricted to specific ML architectures or training schemes; while we employ reservoir computing, the framework is adaptable to other implementations. (4) Capturing Both Short-term and Long-term Behaviors: It predicts both short-term dynamics and long-term statistical properties (climate), making it applicable to a wide range of forecasting tasks. (5) Accuracy and Efficiency with Reservoir Computers: While ML model-agnostic, our implementation leverages the simplicity, efficiency, and low computational costs of RCs, extending the utility of traditional RCs to more complex forecasting scenarios requiring generalization.
Unlike prior work that relies on explicit context indicators/labels to capture the climate of unseen systemsPatel et al. (2021); Kong et al. (2021); Köglmayr and Räth (2024); Panahi and Lai (2024); Kong, Brewer, and Lai (2024), METAFORS requires only short cue signals to construct and initialize forecasting models directly from observations. This label-free learning makes METAFORS versatile: it minimizes the need for domain knowledge and enables learning and prediction across systems governed by distinct functional forms without requiring additional contextual information – a key distinction from generalization schemes that rely on contextual tags.
While METAFORS is not tied to a specific ML architecture, our reservoir computing implementation extends the capabilities of RCs themselves. By mitigating brittleness, data intensity, and warm-up requirements, METAFORS aligns with recent efforts to enhance RC generalization – a central challenge to industrial and scientific applications of RCsYan et al. (2024). Related methods have explored cold starting for RCsGrigoryeva et al. (2024) and LSTMsKemeth et al. (2021), but these approaches focus only on short-term forecasting and assume identical training and test dynamics. Similarly, other meta-learning frameworks often rely on specific architectures, such as convolutional networks for spatiotemporal dataWang, Walters, and Yu (2022) or autoencoders for dimensionality reductionCanaday, Pomerance, and Girvan (2021); Kirchmeyer et al. (2022); Panahi et al. (2025), limiting their flexibility. METAFORS, by contrast, generalizes across diverse systems, cold starts memory-based models, and accommodates a range of architectures for forecasting and signal mapping.
@@
Despite its strengths, our implementation of METAFORS has limitations that warrant further exploration. For example, it requires uniformly sampled sequential data, and adapting the scheme for irregularly sampled or multiscale data would expand its applicability. Additionally, while we demonstrate success in relatively low-dimensional nonlinear systems, scalability to high-dimensional and real-world datasets remains an open question. Future work integrating unsupervised learning techniques, such as autoencoders, could enhance its ability to extract meaningful low-dimensional representations and improve its scalability. Further, while METAFORS’ data-driven modeling approach is an important strength, the integration of METAFORS with hybrid forecasting architectures – where knowledge-based models are coupled with data-driven components – could further enhance its utility. Exploring such integrations could enable METAFORS to address scenarios where partial knowledge of the system dynamics is available, combining the strengths of data-driven and knowledge-based approaches.